On January 2, 2015, Flint Corporation issued $1,750,000 of 10% bonds at 98 due December 31, 2024. Interest on the bonds is payable annually each December 31. The discount on the bonds is also being amortized on a straight-line basis over the 10 years. (Straight-line is not materially different in effect from the preferable “interest method.”)

The bonds are callable at 101 (i.e., at 101% of face amount), and on January 2, 2020, Flint called $1,050,000 face amount of the bonds and redeemed them.

Ignoring income taxes, compute the amount of loss, if any, to be recognized by Flint as a result of retiring the $1,050,000 of bonds in 2020.

Loss on redemption $.

Prepare the journal entry to record the redemption.



Please explain the process in getting to the answer
Answer #1

Issue price of bonds = 1750000*.98 = 1715000

Discount on bonds payable = 1750000-1715000 = 35000

Annually amortization = 35000/10 = 3500

January 2,2020 Carrying value = 1715000+(3500*5) = 1732500

value of 1050000 Bonds = 1732500/1750000*1050000 = 1039500

Loss on redemption = 1050000*1.01-1039500 = 21000

Journal entry

Date account and explanation Debit Credit
Bonds payable 1050000
Loss on redemption on bonds 21000
Discount on bonds payable 10500
Cash 1060500